# On closure operations in the space of subgroups and applications

### v1

@@ Abstract.
We establish some interactions between uniformly recurrent subgroups (URSs) of a group $G$ and cosets topologies $\tau_{\mathcal{N}}$ on $G$ associated to a family $\mathcal{N}$ of normal subgroups of $G$. We show that when $\mathcal{N}$ consists of finite index subgroups of $G$, there is a natural closure operation $\mathcal{H}\mapsto\mathrm{cl}_{\mathcal{N}}(\mathcal{H})$ that associates to a URS $\mathcal{H}$ another URS $\mathrm{cl}_{\mathcal{N}}(\mathcal{H})$, called the $\tau_{\mathcal{N}}$-closure of $\mathcal{H}$. We give a characterization of the URSs $\mathcal{H}$ that are $\tau_{\mathcal{N}}$-closed in terms of stabilizer URSs. This has consequences on arbitrary URSs when $G$ belongs to the class of groups for which every faithful minimal profinite action is topologically free. We also consider the largest amenable URS $\mathcal{A}_{G}$, and prove that for certain coset topologies on $G$, almost all subgroups $H\in\mathcal{A}_{G}$ have the same closure. For groups in which amenability is detected by a set of laws, we deduce a criterion for $\mathcal{A}_{G}$ to be a singleton based on residual properties of $G$.
@@ 1. Introduction
Let $G$ be a group. We denote by $\mathcal{N}_{G}$ the set of normal subgroups of $G$. Let $\mathcal{N}\subseteq\mathcal{N}_{G}$ be a family of normal subgroups of $G$ that is filtering: for every $N_{1},N_{2}\in\mathcal{N}$ there exists $N_{3}\in\mathcal{N}$ such that $N_{3}\leq N_{1}\cap N_{2}$. There is a group topology $\tau_{\mathcal{N}}$ on $G$ associated to $\mathcal{N}$, defined by declaring that the family of cosets $gN$, $g\in G$, $N\in\mathcal{N}$, forms a basis for $\tau_{\mathcal{N}}$. When $\mathcal{N}$ is the family of all finite index normal subgroups of $G$, $\tau_{\mathcal{N}}$ is the profinite topology on $G$. If $p$ is a prime and $\mathcal{N}$ is the family of finite index normal subgroups $N$ of $G$ such that $G/N$ is a $p$-group, $\tau_{\mathcal{N}}$ is the pro-$p$ topology.
@@
In the case of the profinite topology, the notion of pro-$\mathcal{N}$ $G$-space coincides with the classical notion of profinite $G$-space. So in that situation the above proposition says that a URS $\mathcal{H}$ is closed for the profinite topology if and only if $\mathcal{H}$ is the stabilizer URS associated to a minimal profinite action of $G$. Consequences on all URSs can be drawn out of this when $G$ belongs to the class of groups for which, for a faithful minimal compact $G$-space, profinite implies topologically free. See Proposition 4.17. This class of groups includes non-abelian free groups, and more generally any group $G$ admitting an isometric action on a hyperbolic space with unbounded orbits such that the $G$-action on its limit set is faithful. It also includes hereditarily just-infinite groups. Recall that a group $G$ is just-infinite if $G$ is infinite and $G/N$ is finite for every non-trivial normal subgroup $N$, and $G$ is hereditarily just-infinite if every finite index subgroup of $G$ is just-infinite. We call a subgroup $H$ of $G$ quasi-dense in $G$ for the profinite topology if the profinite closure of $H$ has finite index in $G$. For hereditarily just-infinite groups we obtain:
@@
Let $G$ be a hereditarily just-infinite group, and let $\mathcal{H}$ be a non-trivial URS of $G$. Then for every $H\in\mathcal{H}$, $H$ is quasi-dense in $G$ for the profinite topology.
In cases where we know a priori that the group $G$ has the property that the only subgroups that are quasi-dense are the finite index subgroups, we deduce that such a group $G$ admits no continuous URS (a URS is continuous if it is not a finite set). See Corollary 4.20, and the surrounding discussion for context and examples.
@@
The following result provides a criterion for a group to be in $\mathcal{F}$ that is based on residual properties of the group.
@@
The proof from [BKKO17] relies on linear groups technology. Here the argument to deduce Corollary 1 from Theorem 1 uses a reduction to the case of finitely generated groups, and then only appeals to Malcev’s theorem that finitely generated linear groups are residually finite, and the Tits alternative.
@@
The proof of the proposition makes crucial use of the strong proximality of the action of $G$ on $\mathcal{A}_{G}$. The proof of Theorem 1 is easily deduced from the proposition. The additional point is to ensure that the closed normal subgroup $M$ appearing in the conclusion of the proposition remains amenable, and this is where the two assumptions in the theorem are used. We refer to Section 5 for details. Here we only mention that the actual setting in which we prove Theorem 1 does not necessarily require amenable subgroups to be virtually solvable. The assumption that we need is that amenability within subgroups of $G$ can be detected by a set of laws (Definition 5.10), a property that can be thought of as a version of the Tits alternative. See Theorem 5.12 for the more general formulation of the theorem.
Acknowledgements. Thanks are due to Uri Bader and Pierre-Emmanuel Caprace. We can trace back that the possibility of using Proposition 2.1 specifically in the space of subgroups to build a URS starting from another one and a semi-continuous map had been originally brought to our attention by them several years ago.
@@ 2.2. The space of subgroups and URSs
We denote by $\operatorname{Sub}(G)$ the space of subgroups of $G$, equipped with the product topology from $\left\{0,1\right\}^{G}$. It is a compact $G$-space, where $G$ acts by conjugation. If $H\in\operatorname{Sub}(G)$, we denote by $H^{G}$ the $G$-conjugates of $H$, i.e. the $G$-orbit of $H$ in $\operatorname{Sub}(G)$.
@@
If $\mathcal{H}$ is a URS of $G$, we denote by $\mathrm{Env}(\mathcal{H})$ the subgroup generated by all subgroups $H$ in $\mathcal{H}$. The subgroup $\mathrm{Env}(\mathcal{H})$ is normal in $G$, and it is the smallest normal subgroup of $G$ containing some subgroup $H\in\mathcal{H}$.
@@ 2.3. Coset topologies on groups
Let $\mathcal{N}\subseteq\mathcal{N}_{G}$. If $H$ is a subset of $G$, we denote
$\mathrm{cl}_{\mathcal{N}}(H)=\bigcap_{N\in\mathcal{N}}HN.$
@@
    Suppose that $\mathcal{N}$ is filtering. Then for every subset $H$ of $G$, the closure of $H$ with respect to $\tau_{\mathcal{N}}$ is equal to $\mathrm{cl}_{\mathcal{N}}(H)$.
@@ 2.4. Laws
Let $w=w(x_{1},\ldots,x_{k})$ be a word in $k$ letters $x_{1},\ldots,x_{k}$, meaning that $w$ is an element of the free group $F_{k}$ freely generated by $x_{1},\ldots,x_{k}$. Given a group $G$, the word $w$ naturally defines a map $G^{k}\to G$, a $k$-tuple $(g_{1},\ldots,g_{k})$ being mapped to the element $w(g_{1},\ldots,g_{k})$ of $G$ that is obtained by replacing each $x_{i}$ by $g_{i}$. We denote by $\Sigma_{w}(G)\subseteq G^{k}$ the set of $(g_{1},\ldots,g_{k})$ such that $w(g_{1},\ldots,g_{k})=1$. We say that $G$ satisfies the law $w$ if $\Sigma_{w}(G)=G^{k}$.
@@ 3. The $\tau_{\mathcal{N}}$-closure of a URS
Let $\mathcal{H}$ be a closed subset of $\operatorname{Sub}(G)$, and $L$ a subgroup of $G$. For every $\Sigma\subseteq L^{G}$, we write
@@ 4.2. Hereditarily minimal actions
Let $(F_{n})$ be a sequence of non-abelian finite groups. Suppose that for every $n$ there is an abelian subgroup $A_{n}$ of $F_{n}$ such that the only normal subgroup $N$ of $F_{n}$ containing $A_{n}$ is $N=F_{n}$. Let $\mathbb{G}=\prod_{n}F_{n}$, and let $G$ be a countable dense subgroup of $\mathbb{G}$ containing $\bigoplus_{n}F_{n}$. Consider the $G$-action on $X=\prod_{n}F_{n}/A_{n}$. This action is minimal and $G_{x}$ is abelian for every $x\in X$. In particular every $H\in S_{G}(X)$ is abelian. On the other hand $\mathrm{Env}(\mathcal{H})$ contains the normal closure in $G$ of $\bigoplus_{n}A_{n}$. In particular $\mathrm{Env}(\mathcal{H})$ contains $\bigoplus_{n}F_{n}$, and hence $\mathrm{Env}(\mathcal{H})$ is not abelian.
@@
Continuing the previous example, we note that by taking a sequence $(F_{n})$ such that no law $w$ is satisfied by all $F_{n}$ (for instance $F_{n}=\mathrm{Sym}(n)$ for all $n$), we actually obtain an example where $\mathrm{Env}(\mathcal{H})$ satisfies no law at all.
We note that examples of groups as above can be found among finitely generated groups. For instance the groups constructed by B.H. Neumann in [Neu37, Ch. III] satisfy this properties.
@@ 4.3. PIF groups
In this paragraph we focus on the class of groups for which, for a faithful minimal compact $G$-space, profinite implies topologically free.
@@
We say that a subgroup $H$ of a group $G$ is quasi-dense for the profinite topology if $\mathrm{cl}(H)$ is a finite index subgroup of $G$.
@@
Suppose $G$ is HJI, and let $\mathcal{H}$ be a non-trivial URS of $G$. Then for every $H\in\mathcal{H}$, $H$ is quasi-dense in $G$ for the profinite topology.
@@
Following [Cor06], we shall say that a group $G$ has property (PF) if for every subgroup $H$ of $G$, $H$ is quasi-dense in $G$ for the profinite topology only if $H$ has finite index. Let $p$ be a prime number. Following [EJZ13], we say that a group $G$ is weakly $p$-LERF if for every subgroup $H$ of $G$, the closure $\mathrm{cl}_{p}(H)$ of $H$ for the pro-$p$ topology has finite index in $G$ only if $H$ has finite index. Note that applying the definition of weakly $p$-LERF to the trivial subgroup, we see that a just-infinite group that is weakly $p$-LERF is necessarily residually-$p$. Every group that is weakly $p$-LERF has property (PF). Among their striking properties, the finitely generated groups obtained as the outputs of the process carried out in [EJZ13] are HJI and weakly $p$-LERF.
@@
Suppose $G$ is HJI and has property (PF). Then $G$ has no continuous URS.
@@ 5. The Furstenberg URS
Recall that a compact $G$-space $X$ is strongly proximal if the orbit closure of every probability measure on $X$ in the space $\mathrm{Prob}(X)$ contains a Dirac measure. The Furstenberg boundary $\partial_{F}G$ of $G$ is the universal minimal and strongly proximal $G$-space [Gla76]. We denote by $\mathrm{Rad}(G)$ the amenable radical of the group $G$. It coincides with the kernel of the action of $G$ on $\partial_{F}G$.
@@
    If $N\in\mathcal{N}_{G}$ is amenable and if $\operatorname{Sub}_{\geq N}(G)$ is the set of subgroups of $G$ containing $N$, then the natural map $\varphi:\operatorname{Sub}(G/N)\rightarrow\operatorname{Sub}_{\geq N}(G)$ induces a $G$-equivariant homeomorphism between $\mathcal{A}_{G/N}$ and $\mathcal{A}_{G}$.

### v3

@@ Abstract.
We establish some interactions between uniformly recurrent subgroups (URSs) of a group $G$ and cosets topologies $\tau_{\mathcal{N}}$ on $G$ associated to a family $\mathcal{N}$ of normal subgroups of $G$. We show that when $\mathcal{N}$ consists of finite index subgroups of $G$, there is a natural closure operation $\mathcal{H}\mapsto\mathrm{cl}_{\mathcal{N}}(\mathcal{H})$ that associates to a URS $\mathcal{H}$ another URS $\mathrm{cl}_{\mathcal{N}}(\mathcal{H})$, called the $\tau_{\mathcal{N}}$-closure of $\mathcal{H}$. We give a characterization of the URSs $\mathcal{H}$ that are $\tau_{\mathcal{N}}$-closed in terms of stabilizer URSs. This has consequences on arbitrary URSs when $G$ belongs to the class of groups for which every faithful minimal profinite action is topologically free. We also consider the largest amenable URS $\mathcal{A}_{G}$, and prove that for certain coset topologies on $G$, almost all subgroups $H\in\mathcal{A}_{G}$ have the same closure. For groups in which amenability is detected by a set of laws (a property that is variant of the Tits alternative), we deduce a criterion for $\mathcal{A}_{G}$ to be a singleton based on residual properties of $G$.
Keywords: profinite topology and other coset topologies, space of subgroups, uniformly recurrent subgroups, minimal actions on compact spaces, proximal and strongly proximal actions, C*-simplicity.
@@ 1. Introduction
Let $G$ be a group***The main situation we have in mind is when $G$ is countable. Some of our results will require this countability assumption.. We denote by $\mathcal{N}_{G}$ the set of normal subgroups of $G$. Let $\mathcal{N}\subseteq\mathcal{N}_{G}$ be a family of normal subgroups of $G$ that is filtering: for every $N_{1},N_{2}\in\mathcal{N}$ there exists $N_{3}\in\mathcal{N}$ such that $N_{3}\leq N_{1}\cap N_{2}$. There is a group topology $\tau_{\mathcal{N}}$ on $G$ associated to $\mathcal{N}$, defined by declaring that the family of cosets $gN$, $g\in G$, $N\in\mathcal{N}$, forms a basis for $\tau_{\mathcal{N}}$. When $\mathcal{N}$ is the family of all finite index normal subgroups of $G$, $\tau_{\mathcal{N}}$ is the profinite topology on $G$. If $p$ is a prime and $\mathcal{N}$ is the family of finite index normal subgroups $N$ of $G$ such that $G/N$ is a $p$-group, $\tau_{\mathcal{N}}$ is the pro-$p$ topology.
@@
In the case of the profinite topology, the notion of pro-$\mathcal{N}$ $G$-space coincides with the classical notion of profinite $G$-space. So in that situation the above proposition says that a URS $\mathcal{H}$ is closed for the profinite topology if and only if $\mathcal{H}$ is the stabilizer URS associated to a minimal profinite action of $G$. Consequences on all URSs can be drawn out of this when $G$ belongs to the class of groups for which, for a faithful minimal compact $G$-space, profinite implies topologically free. See Proposition 4.17. This class of groups includes non-abelian free groups, and more generally any group $G$ admitting an isometric action on a hyperbolic space with unbounded orbits such that the $G$-action on its limit set is faithful. It also includes hereditarily just-infinite groups. Recall that a group $G$ is just-infinite if $G$ is infinite and $G/N$ is finite for every non-trivial normal subgroup $N$, and $G$ is hereditarily just-infinite if every finite index subgroup of $G$ is just-infinite. We call a subgroup $H$ of $G$ co-finitely dense in $G$ for the profinite topology if the profinite closure of $H$ has finite index in $G$. For hereditarily just-infinite groups we obtain:
@@
Let $G$ be a hereditarily just-infinite group, and let $\mathcal{H}$ be a non-trivial URS of $G$. Then for every $H\in\mathcal{H}$, $H$ is co-finitely dense in $G$ for the profinite topology.
In cases where we know a priori that the group $G$ has the property that the only subgroups that are co-finitely dense are the finite index subgroups, we deduce that such a group $G$ admits no continuous URS (a URS is continuous if it is not a finite set). See Corollary 4.20, and the surrounding discussion for context and examples.
@@
The following result provides a criterion for a group to be in $\mathcal{F}$ that is based on residual properties of the group. If $\mathcal{C}$ is a class of groups, a group $G$ is residually-$\mathcal{C}$ if the intersection of all normal subgroups $N$ such that $G/N\in\mathcal{C}$ is trivial.
@@
Every group satisfying the Tits alternative has the property that every amenable subgroup is virtually solvable. We refer to §5.1 for examples of groups that are known to satisfy the Tits alternative.
@@
The proof from [BKKO17] relies on linear group technology. Here the argument to deduce Corollary 1 from Theorem 1 uses a reduction to the case of finitely generated groups, and then only appeals to Malcev’s theorem that finitely generated linear groups are residually finite, and the Tits alternative [Tit72].
@@
The proof of the proposition makes crucial use of the strong proximality of the action of $G$ on $\mathcal{A}_{G}$. The proof of Theorem 1 is easily deduced from the proposition. The additional point is to ensure that the closed normal subgroup $M$ appearing in the conclusion of the proposition remains amenable, and this is where the two assumptions in the theorem are used. We refer to Section 5 for details. Here we only mention that the actual setting in which we prove Theorem 1 does not necessarily require amenable subgroups to be virtually solvable. The assumption that we need is that amenability within subgroups of $G$ can be detected by a set of laws (Definition 5.10), a property that can be thought of as a version of the Tits alternative (Proposition 5.11). See Theorem 5.13 for the more general formulation of the theorem.
Acknowledgements. Thanks are due to Uri Bader and Pierre-Emmanuel Caprace. We can trace back that the possibility of using Proposition 2.1 specifically in the space of subgroups to build a URS starting from another one and a semi-continuous map had been originally brought to our attention by them several years ago. We also tank a referee who provided various comments that improved the exposition of the paper.
@@ 2.2. The space of subgroups and URSs
We denote by $\operatorname{Sub}(G)$ the space of subgroups of $G$, equipped with the product topology from $\left\{0,1\right\}^{G}$. It is a compact $G$-space, where $G$ acts by conjugation. When $G$ is countable, the space $\operatorname{Sub}(G)$ is second-countable. If $H\in\operatorname{Sub}(G)$, we denote by $H^{G}$ the $G$-conjugates of $H$, i.e. the $G$-orbit of $H$ in $\operatorname{Sub}(G)$.
@@
If $\mathcal{H}$ is a URS of $G$, we denote by $\mathrm{Env}(\mathcal{H})$ the subgroup generated by all subgroups $H$ in $\mathcal{H}$, called the envelope of $\mathcal{H}$. The subgroup $\mathrm{Env}(\mathcal{H})$ is normal in $G$, and it is the smallest normal subgroup of $G$ containing some subgroup $H\in\mathcal{H}$.
@@ 2.3. Coset topologies on groups
Let $\mathcal{N}\subseteq\mathcal{N}_{G}$. If $E$ is a subset of $G$, we denote
$\mathrm{cl}_{\mathcal{N}}(E)=\bigcap_{N\in\mathcal{N}}EN.$
@@
    Suppose that $\mathcal{N}$ is filtering. Then for every subset $E$ of $G$, the closure of $E$ with respect to $\tau_{\mathcal{N}}$ is equal to $\mathrm{cl}_{\mathcal{N}}(E)$.
@@ 2.4. Laws
Let $w=w(x_{1},\ldots,x_{k})$ be a word in $k$ letters $x_{1},\ldots,x_{k}$, meaning that $w$ is an element of the free group $F_{k}$ freely generated by $x_{1},\ldots,x_{k}$. We assume $w$ is non-trivial. Given a group $G$, the word $w$ naturally defines a map $G^{k}\to G$, a $k$-tuple $(g_{1},\ldots,g_{k})$ being mapped to the element $w(g_{1},\ldots,g_{k})$ of $G$ that is obtained by replacing each $x_{i}$ by $g_{i}$. We denote by $\Sigma_{w}(G)\subseteq G^{k}$ the set of $(g_{1},\ldots,g_{k})$ such that $w(g_{1},\ldots,g_{k})=1$. We say that $G$ satisfies the law $w$ if $\Sigma_{w}(G)=G^{k}$.
Example 2.9.
Abelian groups are the groups that satisfy the law $w_{1}=[x_{1},x_{2}]=x_{1}x_{2}x_{1}^{-1}x_{2}^{-1}$. More generally let $(w_{\ell})_{\ell\geq 1}$ be the sequence of laws defined inductively by the rule: if $w_{\ell}$ is already defined and involves $2^{\ell}$ variables, then
$w_{\ell+1}(x_{1},\ldots,x_{2^{\ell+1}})=[w_{\ell}(x_{1},\ldots,x_{2^{\ell}}),w_{\ell}(x_{2^{\ell}+1},\ldots,x_{2^{\ell+1}})].$
Then a group $G$ is solvable of length at most $\ell$ if and only if $G$ satisfies the law $w_{\ell}$.
@@ 3. The $\tau_{\mathcal{N}}$-closure of a URS
Let $\mathcal{H}$ be a closed subset of $\operatorname{Sub}(G)$, and $L$ a subgroup of $G$. Recall that $L^{G}$ is the set of $G$-conjugates of $L$. For every $\Sigma\subseteq L^{G}$, we write
@@
The relevance of this definition is indicated by the following two lemmas.
@@ 4.2. Hereditarily minimal actions
Recall that every minimal and proximal $G$-space is hereditarily minimal [Gla76, Lemma 3.2].
@@
Let $(F_{n})$ be a sequence of non-abelian finite groups. Suppose that for every $n$ there is an abelian subgroup $E_{n}$ of $F_{n}$ such that the only normal subgroup $N$ of $F_{n}$ containing $E_{n}$ is $N=F_{n}$. Let $\mathbb{G}=\prod_{n}F_{n}$, and let $G$ be a countable dense subgroup of $\mathbb{G}$ containing $\bigoplus_{n}F_{n}$. Consider the $G$-action on $X=\prod_{n}F_{n}/E_{n}$. This action is minimal and $G_{x}$ is abelian for every $x\in X$. In particular every $H\in\mathcal{H}:=S_{G}(X)$ is abelian. On the other hand $\mathrm{Env}(\mathcal{H})$ contains the normal closure in $G$ of $\bigoplus_{n}E_{n}$. In particular $\mathrm{Env}(\mathcal{H})$ contains $\bigoplus_{n}F_{n}$, and hence $\mathrm{Env}(\mathcal{H})$ is not abelian.
Examples as above can be found among finitely generated groups. For instance the groups constructed by B.H. Neumann in [Neu37, Ch. III] satisfy these properties. Given a sequence $\mathbf{u}=(u_{n})_{n}$ of odd integers greater than $5$, B.H. Neumann constructs a subgroup $G_{\mathbf{u}}$ of the product $\prod_{n}\mathrm{Alt}(u_{n})$. The subgroup $G_{\mathbf{u}}$ is generated by two elements, and $G_{\mathbf{u}}$ contains $\bigoplus_{n}\mathrm{Alt}(u_{n})$ [Neu37, Ch. III (16)]. Since $\mathrm{Alt}(u_{n})$ is simple, $G_{\mathbf{u}}$ therefore falls into the above setting with $F_{n}=\mathrm{Alt}(u_{n})$ and $E_{n}$ any non-trivial abelian subgroup of $\mathrm{Alt}(u_{n})$.
@@
We note that in the above examples $G_{\mathbf{u}}$, we actually have an abelian URS $\mathcal{H}$ for which $\mathrm{Env}(\mathcal{H})$ satisfies no law at all (since a given non-trivial law cannot be satisfied by all finite groups, and hence neither by arbitrary large finite alternating groups).
@@ 4.3. PIF groups
In this subsection we focus on the class of groups for which, for a faithful minimal compact $G$-space, profinite implies topologically free.
@@
We say that a subgroup $H$ of a group $G$ is co-finitely dense for the profinite topology if $\mathrm{cl}(H)$ is a finite index subgroup of $G$.
@@
Suppose $G$ is hereditarily just-infinite, and let $\mathcal{H}$ be a non-trivial URS of $G$. Then for every $H\in\mathcal{H}$, $H$ is co-finitely dense in $G$ for the profinite topology.
@@
Following [Cor06], we shall say that a group $G$ has property (PF) if for every subgroup $H$ of $G$, $H$ is co-finitely dense in $G$ for the profinite topology only if $H$ has finite index. Let $p$ be a prime number. Following [EJZ13], we say that a group $G$ is weakly $p$-LERF if for every subgroup $H$ of $G$, the closure $\mathrm{cl}_{p}(H)$ of $H$ for the pro-$p$ topology has finite index in $G$ only if $H$ has finite index. Note that applying the definition of weakly $p$-LERF to the trivial subgroup, we see that a just-infinite group that is weakly $p$-LERF is necessarily residually-$p$. Every group that is weakly $p$-LERF has property (PF). Among their striking properties, the finitely generated groups obtained as the outputs of the process carried out in [EJZ13] are HJI and weakly $p$-LERF.
@@
Suppose $G$ is hereditarily just-infinite and has property (PF). Then $G$ has no continuous URS.
@@ 5. The Furstenberg URS
Recall that a compact $G$-space $X$ is strongly proximal if the orbit closure of every probability measure on $X$ in the space $\mathrm{Prob}(X)$ contains a Dirac measure. The Furstenberg boundary $\partial_{F}G$ of $G$ is the universal minimal and strongly proximal $G$-space [Fur73], [Gla76]. We denote by $\mathrm{Rad}(G)$ the amenable radical of the group $G$. It coincides with the kernel of the action of $G$ on $\partial_{F}G$ (a result that holds more generally for arbitrary locally compact groups [Fur03]).
@@
    If $N$ is an amenable normal subgroup of $G$, and if $\operatorname{Sub}_{\geq N}(G)$ is the set of subgroups of $G$ containing $N$, then the natural map $\varphi:\operatorname{Sub}(G/N)\rightarrow\operatorname{Sub}_{\geq N}(G)$ induces a $G$-equivariant homeomorphism between $\mathcal{A}_{G/N}$ and $\mathcal{A}_{G}$.